OpenStudy (minecraftqueen):

What is the slope of the line passing through the points (1, –5) and (4, 1)? A.2 B. -4/5 C. 4/5 D.-2

OpenStudy (sleepyjess):

Hello! Do you know the slope formula?

OpenStudy (minecraftqueen):

x1-x2 y1-y2 ?

OpenStudy (sleepyjess):

Close, \(m = \dfrac{y_2 - y_1}{x_2-x_1}\)

OpenStudy (minecraftqueen):

-5-1/1-4 ?

OpenStudy (sleepyjess):

Almost! It would be 4 - 1 on the bottom

OpenStudy (minecraftqueen):

(1 /-5)/(4-1) = 1/20

OpenStudy (sleepyjess):

Where'd you get all that from? Remember you had it right the first time, except you needed to flip the bottom

OpenStudy (sleepyjess):

\(\dfrac{-5-1}{4-1}\)

OpenStudy (minecraftqueen):

-10

OpenStudy (sleepyjess):

-10? I'm not sure what you're doing

OpenStudy (minecraftqueen):

sorry im kinda in a hurry ill just quess

OpenStudy (sleepyjess):

What is -5 - 1?

OpenStudy (minecraftqueen):

i gtg baii

OpenStudy (sleepyjess):

....

OpenStudy (minecraftqueen):

@sleepyjess it was -2 some one told me -_- but gtg bai

OpenStudy (sleepyjess):

But what is learned if someone just gives out the answer?

OpenStudy (minecraftqueen):

@sleepyjess didnt have time to tho